# Anisotropic Green Coordinates

### v1

@@ Abstract.
We live in a world filled with anisotropy, a ubiquitous characteristic of both natural and engineered systems. In this study, we concentrate on space deformation and introduce anisotropic Green coordinates, which provide versatile effects for cage-based and variational deformations in both two and three dimensions. The anisotropic Green coordinates are derived from the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a symmetric positive definite matrix. This equation belongs to the class of constant-coefficient second-order elliptic equations, exhibiting properties analogous to the Laplacian equation but incorporating the matrix $\mathbf{A}$ to characterize anisotropic behavior. Based on this equation, we establish the boundary integral formulation, which is subsequently discretized to derive anisotropic Green coordinates defined on the vertices and normals of oriented simplicial cages. The deformation satisfies basic properties such as linear reproduction and translation invariance, and has closed-form expressions for both 2D and 3D scenarios. We also offer intuitive geometric interpretations of this method. Furthermore, our approach computes the gradient and Hessian of the deformation coordinates and employs the local-global optimization framework to facilitate variational shape deformation, enabling flexible shape manipulation while achieving as-rigid-as-possible shape deformation. Experimental results demonstrate that anisotropic Green coordinates offer versatile and diverse deformation options, providing artists with enhanced flexibility and introducing a novel perspective on spatial deformation.
@@ 1. Introduction
Space deformation is a basic problem in geometric modeling. Instead of directly modifying the object surface, space deformation treats the surface as an embedding in the ambient space and deforms the space through deformation handles (2009VariationalGC). Cage-based deformation is a typical example in which the embedded object deforms according to the manipulation of the source cage to the target cage. For each point $\eta$ within the cage, a series of functions $\phi_{i}(\eta)$ is constructed for each cage vertex $\mathbf{v}_{i}$. These functions satisfy the partition of unity property, $\sum_{i=1}^{n}{\phi_{i}(\eta)}=1$, and the linear reproduction property $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}$. The functions $\phi_{i}(\eta)$ are always called the Generalized Barycentric Coordinates (GBC) (2015ReviewBarycentric; 2017HormannBarycentric).
@@
GBC is usually interpolatory and expresses the point inside the cage as an affine sum of the cage vertices. Although it demonstrates good consistency with the target cage, it introduces obvious shearing artifacts, particularly when the cage undergoes affine transformations (2008Green). In practice, it is also desirable for the deformed object to maintain a certain degree of conformality with the original object. Harmonic functions often provide a solution to this issue, as harmonicity is inherently associated with shape preservation. Based on the Green’s third identity of harmonic functions, 2008Green propose Green coordinates, which take into account both the cage vertices $\mathbf{v}_{i}$ and the cage normals $\mathbf{n}_{j}$. Green coordinates satisfy $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}+\sum_{j=1}^{m}\psi_{j}(\eta)\mathbf{n}_{j}$ for $\eta$ inside the cage. These coordinates enable shape-preserving deformations, characterized by angle-preserving in 2D and quasi-conformality in 3D. However, conformality is a strict and absolute condition. While it provides desirable properties, it also restricts the deformation diversity of Green coordinates once the source and target cages are fixed. In practical applications and artistic design, there is often a demand for more flexible and versatile deformation styles, which can be achieved through adjustments to various parameters. In the real world, materials such as bones, wood, and fibers exhibit anisotropic structures, leading users to desire deformations that are either more “flexible” or more “rigid” along specific directions.
@@
Based on previous observations, our work proposes to introduce the anisotropic property into the traditional Green coordinates. The Laplacian operator is the composition of divergence and gradient, i.e., $\Delta=\nabla\cdot\nabla$. Therefore, we can introduce anisotropy within the divergence operator and consider the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a positive-definite matrix. The matrix $\mathbf{A}$ introduces varying scales in different directions, thereby imparting anisotropic properties in the formulation. Furthermore, the anisotropic Laplacian equation also gives rise to corresponding boundary integral formulation and leads to the anisotropic identity equations, which form the foundation for constructing the anisotropic Green coordinates. We show that anisotropic Green coordinates satisfy linear reproduction and translation invariance properties, and produce versatile deformation results with varying control matrices $\mathbf{A}$. Additionally, we provide closed-form expressions for both 2D and 3D cases and offer intuitive geometric interpretations of the method. Through extensive experiments, we validate that anisotropic Green coordinates provide artists with a broader range of deformation options and enhance the diversity of deformations choices. Moreover, appropriate selection of the matrix $\mathbf{A}$ enables desired effects, such as tighter adherence to the deformed cage or reduced isotropic and area distortion.
@@
To further demonstrate the broad applicability of the closed-form anisotropic Green coordinates, we also compute their gradients and Hessians and perform variational shape deformation (2009VariationalGC), which minimizes the energy functional of the deformation map. This approach enables flexible control through positional constraints and supports the as-rigid-as-possible deformation strategy. The incorporation of anisotropy further extends the applicability of the variational shape deformation framework to a wider range of scenarios.
@@ 2.1. Cage-based Deformation
Cage-based deformation is a fundamental topic in geometry processing, where users manipulate a coarse control structure to drive the deformation of complex objects within the enclosed space. Two critical factors in this problem are the geometry of the cage structure and the deformation coordinates defined on the cage. The most commonly used cage structure is the oriented simplicial surface, such as closed polygonal chains in 2D and triangle meshes in 3D (2005MVC; 2008Green; 2009VariationalGC; 2012Biharmonic; 2019GeneralBarycentric; 2024Biharmonic3D). It is the first-order approximation of co-dimensional one surfaces. Other methods utilize high-order surfaces, such as polynomial or Bézier surfaces, as the cage structure (2013CubicMVC; 2023PolyGreen; 2024GCpatych; 2024PolyHighOrder; 2024PolyCauchy; 2025VariationalBiharmonic; 2025BezierCage; 2025Polynomial3D). Such structures enable more precise and flexible control over curved and bending deformations.
@@
The cage structure mainly influences the flexibility of the deformation; however, the deformation effect is determined by the specific deformation coordinates established on the cage. The most widely utilized methods are generalized barycentric coordinates (GBC), which formulate interpolatory coordinates as a partition of unity and are capable of accurately representing any point $\eta$ within the cage. As pioneering works, 2003MVCPoly, 2005MVCPoly, and 2005MVC construct closed-form mean value coordinates by projecting the enclosing cage onto a sphere centered at $\eta$ and calculating the contribution of each cage vertex based on this spherical projection. There are also other techniques for constructing GBC, such as using the Laplacian equation (2007Harmonic), the Poisson integral formula (2013Poisson), probability theory (2008MaxEntropy; 2023MaximumLikelihood), or stochastic sampling and Monte Carlo integration (2024Stochastic). Additionally, transfinite coordinates (2006Transfinite; 2009TransfiniteMVC; 2025TransfiniteArbitary) extend the concept of traditional discrete barycentric coordinates to continuous boundaries, offering a general theoretical framework.
@@
In addition to barycentric coordinates, some other deformation coordinates consider both the cage vertices and cage normals, and therefore achieve shape-preserving deformations. Shape preservation means that the deformed shape maintains a certain degree of conformality with the original shape and avoids large shearing artifacts. The typical methods are the Green coordinates (2008Green; 2009DriveGreen; 2022QuadGreen; 2025Polynomial3D; 2025BezierCage), which utilize Green’s third identity to establish the deformation formulation and naturally satisfy linear reproduction and transformation invariance. Other techniques involve Cauchy coordinates (2009CBC; 2024PolyCauchy; 2025CauchyClose), which is demonstrated to be geometrically equivalent to Green coordinates in the planar domain. However, the deformation effects of Green coordinates and Cauchy coordinates are fixed. Specifically, once the source and target cages are determined, the deformation is uniquely defined. To address this limitation, one approach is to solve the biharmonic Dirichlet problem, which incorporates gradient variations into the formulation (2012Biharmonic; 2024Biharmonic3D; 2025PolyBiharmonic). Alternatively, the variational deformation technique can be employed to optimize the deformation map, as discussed in the following subsection.
@@ 2.2. Variational Shape Deformation
Variational methods are widely used in formulating partial differential equations for practical problems, such as the minimal surface equation (1993MinimalSurface). They address the most fundamental problem in functional analysis of finding a function $f$ that minimizes the energy functional $E(f)$.  2009VariationalGC introduce the variational harmonic map in space deformation, which defines $E(f)$ based on Green coordinates, enabling deformation control with position constraints and as-rigid-as-possible properties. The optimization problem is solved through a local-global minimization strategy (2007ARAP; 2008LocalGlobal). This method requires the computation of the Jacobian and Hessian of the Green coordinate. Additionally, some subsequent studies extend variational shape deformation techniques using similar strategies across different coordinate systems and high-order cage structures (2012Biharmonic; 2024Biharmonic3D; 2025CauchyClose; 2025VariationalBiharmonic; 2025Polynomial3D). Furthermore, 2023VBC introduce variational barycentric coordinates, a neural field-based framework that optimizes generalized barycentric coordinates while minimizing the total variation.
@@ 2.3. Anisotropy in Geometry Processing
At the mathematical fundamental level, 2007Upscaling establish a metric-based upscaling framework for elliptic PDEs of the form $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))$ and rigorously analyze its mathematical properties within the Sobolev space.  2009InhomogeneousElastic construct anisotropic elasticity tensors on coarse meshes through harmonic homogenization to preserve fine-scale directional properties. 2019AnisotropicElasticity focus on anisotropic hyperelasticity and propose an anisotropic strain invariant for the inversion-safe energy model, complemented by a rehabilitation technique to address degenerate elements. 2024AnisotropicFiniteElement derive closed-form expressions for anisotropic invariants, facilitating the stable simulation of directionally dependent materials.
Anisotropy is also well introduced in mesh processing. 2022GoGreen expand Green’s functions by leveraging the anisotropic characteristics of general linear elastic materials, which allows for real-time control of material deformation without volumetric discretization. 2018AnisotropicQuasiConformal propose generating planar anisotropic meshes for a given metric tensor and establishing a quasi-conformal mapping to transform the anisotropic metric into a Euclidean metric. 2013ParticleSurfaceMeshing embed the anisotropic Riemannian metrics into higher-dimensional isotropic spaces, achieving efficient particle-based anisotropic meshing. 2025AnisotropicGauss derive the anisotropic Gauss formulation and utilize this boundary integral formulation for point cloud normal orientation and surface reconstruction. This work focuses on geometry topics that are distinct from our method and only considers the anisotropic equations in diagonal forms.
@@ 3. Anisotropic Green Coordinates
In this section, we present a detailed derivation of the anisotropic Green coordinates. We begin by introducing the boundary integral formulation of the anisotropic Laplacian equation. It is not trivial that, despite the introduction of anisotropy, the equation can still be expressed using boundary integrals to determine the coordinates of any point within the cage. Subsequently, we discretize this integral equation to obtain anisotropic Green coordinates that exhibit linear reproduction and translation invariance. Finally, we analyze their properties.
@@ 3.1. Anisotropic Laplacian Equation
where $\mathbf{A}$ is a $d\times d$ positive-definite matrix with constant coefficients. This is the simplification of the second-order elliptic operator $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))+\mathbf{b}(\mathbf{x})\cdot\nabla u(\mathbf{x})+cu(\mathbf{x})$ (1064EvansPDE), which shares similar properties with harmonic functions. When $\mathbf{A}$ is not a constant matrix, the properties of its solutions are usually analyzed in the Sobolev space. Deriving explicit solutions is often challenging. Thus, we focus on the scenario where $\mathbf{A}$ is a constant-coefficient matrix, and both the convection-enhancing term $\mathbf{b}(\mathbf{x})$ and the first-order term $c$ are set to zero. It has been mathematically established that every differential operator $L$ with constant coefficients has a fundamental solution (1995PDEbook).
We first present a preliminary of the square roots of the positive definite matrix. We know that every positive definite matrix $\mathbf{A}$ can be expressed as $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$, where
@@
is a $d$-dimensional diagonal matrix, and $\mathbf{P}$ is an orthogonal matrix satisfying $\mathbf{P}^{-1}=\mathbf{P}^{\top}$ (2003Strang). By expressing
@@
we can define the square root of $\mathbf{A}$ as $\mathbf{A}^{1/2}=\mathbf{P}\Lambda^{1/2}\mathbf{P}^{\top}$. It is then straightforward to verify that $(\mathbf{A}^{1/2})^{2}=\mathbf{A}$.
Next, we will present some theorems related to the anisotropic Laplacian equation. While we believe these properties have been investigated in prior fundamental mathematical research, as they are closely related to the Green’s function (2004GeneralizedGreen), we have not identified specific literature in the fields of cage or variational deformation that comprehensively addresses these results. Therefore, we provide a detailed discussion of them in the following.
@@
where $\delta(\mathbf{x})$ is the Dirac delta distribution satisfying $\delta(\mathbf{x})=0$ for $\mathbf{x}\neq\mathbf{0}$ and $\int_{\mathbb{R}^{d}}{\delta(\mathbf{x})=1}$, and $\omega_{d}=2\pi/\Gamma(n/2)$ represents the surface area of the unit sphere in $\mathbb{R}^{d}$.
@@
Let $\mathbf{x}=\mathbf{A}^{-1/2}\xi$, $\mathbf{y}=\mathbf{A}^{-1/2}\eta$. According to lemma A.1 in the supplementary material, and the fact that $\mathbf{A}^{-1/2}$ is symmetric and invertible, we have $\nabla_{\xi}f=\mathbf{A}^{-1/2}\nabla_{\mathbf{x}}f$ and $\nabla_{\xi}\cdot\mathbf{F}=\nabla_{\mathbf{x}}\cdot(\mathbf{A}^{-1/2}\mathbf{F})$, where $f$ and $\mathbf{F}$ are scalar and vector value functions. Then,
@@
For the rightmost equality, refer to Lemma A.2 in the appendix, together with the fact that $\det{\mathbf{A}^{1/2}}=|\det{\Lambda^{1/2}|}=\sqrt{\det{\mathbf{A}}}$. Therefore, we conclude that $\sqrt{\det{\mathbf{A}}}\ G_{\mathbf{A}}(\mathbf{x}-\mathbf{y})\$is the fundamental solution $G(\mathbf{u})$ of $d$-dimensional isotropic Laplacian equation, which is given by
@@
Moreover, we can directly compute $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)$, which is the gradient of the anisotropic fundamental solution, and it is expressed as follows:
(8) $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)=\begin{cases}\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-1}\mathbf{A}^{-1}(\xi-\eta)},&d=2,\\ \frac{1}{\omega_{d}\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-\frac{d}{2}}\mathbf{A}^{-1}(\xi-\eta)},&d\geq 3.\end{cases}$

### v2

@@ Abstract.
We live in a world filled with anisotropy, a ubiquitous characteristic of both natural and engineered systems. In this study, we concentrate on space deformation and introduce anisotropic Green coordinates, which provide versatile effects for cage-based and variational deformations in both two and three dimensions. The anisotropic Green coordinates are derived from the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a symmetric positive definite matrix. This equation belongs to the class of constant-coefficient second-order elliptic equations, exhibiting properties analogous to the Laplacian equation but incorporating the matrix $\mathbf{A}$ to characterize anisotropic behavior. Based on this equation, we establish the boundary integral formulation, which is subsequently discretized to derive anisotropic Green coordinates defined on the vertices and normals of oriented simplicial cages. Our method satisfies basic properties such as linear reproduction and translation invariance, and possesses closed-form expressions for both 2D and 3D scenarios. We also give an intuitive geometric interpretation of the approach, demonstrating that our method generates a quasi-conformal mapping. Furthermore, we derive the gradients and Hessians of the deformation coordinates and employ the local-global optimization framework to facilitate variational shape deformation, enabling flexible shape manipulation while achieving as-rigid-as-possible shape deformation. Experimental results demonstrate that anisotropic Green coordinates offer versatile and diverse deformation options, providing artists with enhanced flexibility and introducing a novel perspective on spatial deformation.
@@ 1. Introduction
Space deformation is a basic problem in geometric modeling. Instead of directly modifying the object surface, space deformation treats the surface as an embedding in the ambient space and deforms the space through deformation handles (Ben-Chen et al., 2009). Cage-based deformation is a typical example in which the embedded object deforms according to the manipulation of the cage (Ströter et al., 2024). For each point $\eta$ within the cage, a series of functions $\phi_{i}(\eta)$ is constructed for each cage vertex $\mathbf{v}_{i}$. These functions satisfy the partition of unity property, $\sum_{i=1}^{n}{\phi_{i}(\eta)}=1$, and the linear reproduction property $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}$. The functions $\phi_{i}(\eta)$ are commonly called the Generalized Barycentric Coordinates (GBC) (Floater, 2015; Hormann and Sukumar, 2017).
@@
GBC is usually interpolatory and expresses the point inside the cage as an affine sum of the cage vertices. Although it demonstrates good consistency with the target cage, it introduces obvious shearing artifacts, particularly when the cage undergoes affine transformations (Lipman et al., 2008). In practice, it is also desirable for the deformed object to maintain a certain degree of conformality with the original object. Based on the Green’s third identity of harmonic functions, Lipman et al. (2008) propose Green coordinates, which take into account both the cage vertices $\mathbf{v}_{i}$ and the cage normals $\mathbf{n}_{j}$. Green coordinates satisfy $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}+\sum_{j=1}^{m}\psi_{j}(\eta)\mathbf{n}_{j}$ for $\eta$ inside the cage. These coordinates enable shape-preserving deformations, characterized by angle-preserving in 2D and quasi-conformality in 3D. However, conformality is a strict and absolute condition, which also restricts the deformation diversity of Green coordinates once the source and target cages are fixed. In practical applications and artistic design, there is often a demand for more flexible and versatile deformation styles, which can be achieved through adjusting various parameters. In the real world, materials such as bones, wood, and fibers exhibit anisotropic structures, prompting users to seek deformations that are more “flexible” or more “rigid” along specific directions.
@@
Based on previous observations, we propose to incorporate the anisotropic property into the traditional Green coordinates. The Laplacian operator is the composition of divergence and gradient, i.e., $\Delta=\nabla\cdot\nabla$. Therefore, we can introduce anisotropy within the divergence operator and consider the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a positive-definite matrix. The matrix $\mathbf{A}$ introduces varying scales in different directions, thereby incorporating anisotropic properties in the formulation. Furthermore, the anisotropic Laplacian equation also gives rise to corresponding boundary integral formulation and leads to the anisotropic Green’s identity, which form the foundation for constructing the anisotropic Green coordinates. We show that anisotropic Green coordinates satisfy linear reproduction and translation invariance properties, and produce versatile deformation results with varying control matrices $\mathbf{A}$. Additionally, we provide closed-form expressions for both 2D and 3D cases and offer intuitive geometric interpretations of the method. We also relate the distortion of the deformation mapping to the condition number of $\mathbf{A}$, and mathematically demonstrate that anisotropic Green coordinates exhibit quasi-conformal properties. Through extensive experiments, we validate that anisotropic Green coordinates provide artists with a broader range of deformation options and enhance the diversity of deformations choices. Moreover, appropriate selection of the matrix $\mathbf{A}$ enables desired effects, such as tighter adherence to the deformed cage or reduced isotropic and area distortion.
@@
To further demonstrate the broad applicability of the closed-form anisotropic Green coordinates, we also derive their gradients and Hessians and perform variational shape deformation (Ben-Chen et al., 2009), which minimizes energy functional of the deformation map. This approach enables flexible control through space constraints and supports the as-rigid-as-possible deformation strategy. The incorporation of anisotropy further extends the applicability of the variational shape deformation framework to a wider range of scenarios.
@@ 2.1. Cage-based Deformation
Cage-based deformation is a fundamental topic in geometry processing, where users manipulate a coarse control structure to drive the deformation of complex objects within the enclosed space (Ströter et al., 2024). Two critical factors in this problem are the geometry of the cage structure and the deformation coordinates defined on the cage. The most commonly used cage structure is the oriented simplicial surface, such as closed polygonal chains in 2D and triangle meshes in 3D (Ju et al., 2005; Lipman et al., 2008; Ben-Chen et al., 2009; Weber et al., 2012; Yan and Schaefer, 2019; Thiery et al., 2024). It is the first-order approximation of co-dimensional one surfaces. Other methods utilize high-order surfaces, such as polynomial or Bézier surfaces, as the cage structure (Li et al., 2013; Michel and Thiery, 2023; Qin et al., 2024; Liu et al., 2024; Lin and Chen, 2024; Michel et al., 2025; Xiao and Chen, 2025; Wu et al., 2025). Such structures enable more precise and flexible control over curved and bending deformations.
@@
While the cage structure can influence the flexibility, the deformation effect is predominantly determined by the specific deformation coordinates established on the cage. The most widely utilized methods are generalized barycentric coordinates (GBC), which formulate interpolatory coordinates as a partition of unity and can accurately represent an arbitrary point $\eta$ within the cage. As pioneering works, Floater (2003), Floater et al. (2005), and Ju et al. (2005) construct closed-form mean value coordinates by projecting the enclosing cage onto a sphere centered at $\eta$ and calculating the contribution of each cage vertex based on this spherical projection. There are also other techniques for constructing GBC, such as using the Laplacian equation (Joshi et al., 2007), Poisson integral formula (Li and Hu, 2013), probability theory (Hormann and Sukumar, 2008; Chang et al., 2023), or stochastic sampling and Monte Carlo integration (de Goes and Desbrun, 2024). Additionally, transfinite coordinates (Belyaev, 2006; Dyken and Floater, 2009; Chang and Hormann, 2025) extend the concept of traditional discrete barycentric coordinates to continuous boundaries, offering a general theoretical framework.
@@
In addition to barycentric coordinates, some other deformation coordinates consider both the cage vertices and cage normals, and therefore achieve shape-preserving deformations. Shape preservation means that the deformed shape maintains a certain degree of conformality with the original shape and avoids large shearing artifacts. The typical methods are the Green coordinates (Lipman et al., 2008; Lipman and Levin, 2009; Thiery and Boubekeur, 2022; Wu et al., 2025; Xiao and Chen, 2025), which utilize Green’s third identity to establish the deformation formulation and naturally satisfy linear reproduction and transformation invariance. Other techniques involve Cauchy coordinates (Weber et al., 2009; Lin and Chen, 2024; Liu et al., 2025b), which is demonstrated to be geometrically equivalent to Green coordinates in the planar domain. However, the deformation effects of Green coordinates and Cauchy coordinates are fixed. Specifically, once the source and target cages are determined, the deformation is uniquely defined. To address this limitation, some approach solve the biharmonic Dirichlet problem, which incorporates gradient variations into the formulation (Weber et al., 2012; Thiery et al., 2024; Liu et al., 2025a). Alternatively, Somigliana coordinates  (Chen et al., 2023) generalize Green coordinates based on the Navier–Cauchy equation of linear elasticity. This approach enables physically plausible volume control while preserving similarity invariance. However, it lacks a closed-form solution and requires a CUDA-dependent implementation. Another option is to use variational deformation techniques to optimize the deformation map, as discussed in the following subsection.
@@ 2.2. Variational Shape Deformation
Variational methods are widely used in formulating partial differential equations for practical problems, such as the minimal surface equation (Pinkall and Polthier, 1993). They address the most fundamental problem in functional analysis of finding a function $f$ that minimizes the energy functional $E(f)$.  Ben-Chen et al. (2009) introduce the variational harmonic map in space deformation, which defines $E(f)$ based on Green coordinates, enabling deformation control with space constraints and as-rigid-as-possible properties. The optimization problem is solved through a local-global minimization strategy (Sorkine and Alexa, 2007; Liu et al., 2008), which requires computing the Jacobian and Hessian of Green coordinates. Subsequent research has extended this technique by incorporating various coordinate systems or employing higher-order cage structures (Weber et al., 2012; Thiery et al., 2024; Liu et al., 2025b; Michel et al., 2025; Wu et al., 2025). Additionally, Dodik et al. (2023) introduce variational barycentric coordinates, a neural field-based framework that optimizes generalized barycentric coordinates by minimizing their total variation.
@@ 2.3. Anisotropy in Geometry Processing
At the mathematical fundamental level, Owhadi and Zhang (2007) establish a metric-based upscaling framework for elliptic PDEs of the form $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))$ and rigorously analyze its mathematical properties within the Sobolev space.  Kharevych et al. (2009) construct anisotropic elasticity tensors on coarse meshes through harmonic homogenization to preserve fine-scale directional properties. Kim et al. (2019) focus on anisotropic hyperelasticity and propose an anisotropic strain invariant for the inversion-safe energy model, complemented by a rehabilitation technique to address degenerate elements. Lin et al. (2024) derive the anisotropic as-rigid-as-possible energy for elastic materials and provide a robust formulation suitable for finite element simulations.
Anisotropy is also well introduced in mesh processing. Chen and Desbrun (2022) expand Green’s functions by leveraging the anisotropic characteristics of general linear elastic materials, which allows for real-time control of material deformation without volumetric discretization. Ren et al. (2018) propose generating planar anisotropic meshes for a given metric tensor and establishing a quasi-conformal mapping to transform the anisotropic metric into a Euclidean metric. Zhong et al. (2013) embed the anisotropic Riemannian metrics into higher-dimensional isotropic spaces, achieving efficient particle-based anisotropic meshing. Ma et al. (2025) derive the anisotropic Gauss formulation and utilize this boundary integral formulation for point cloud normal orientation and surface reconstruction. This work focuses on geometry topics that are distinct from our method and only considers the anisotropic equations in diagonal forms.
@@ 3. Anisotropic Green Coordinates
In this section, we present a detailed derivation of the anisotropic Green coordinates. We begin by introducing the boundary integral formulation of the anisotropic Laplacian equation. It is non-trivial that, despite the introduction of anisotropy, we can still use boundary integrals to determine the coordinates of any point within the cage. Subsequently, we discretize this identity to obtain anisotropic Green coordinates that exhibit linear reproduction and translation invariance. Then, we analyze their properties.
@@ 3.1. Anisotropic Laplacian Equation
where $\mathbf{A}$ is a $d\times d$ positive-definite matrix with constant coefficients. This equation is a special case of the second-order elliptic operator $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))+\mathbf{b}(\mathbf{x})\cdot\nabla u(\mathbf{x})+c(\mathbf{x})u(\mathbf{x})$, and shares many similar properties with the classical Laplacian operator. When $\mathbf{A}$ is not constant, properties of solutions are generally studied in Sobolev spaces (Evans and Lawrence, 1964), and obtaining explicit solutions is often challenging. Thus, we focus on the scenario where $\mathbf{A}$ is a constant-coefficient matrix, and both the convection-enhancing term $\mathbf{b}(\mathbf{x})$ and the first-order term $c(\mathbf{x})$ are set to zero. It has been mathematically established that every differential operator $\mathcal{L}$ with constant coefficients has a fundamental solution (Folland, 1995).
We first introduce the concept of the square root of a positive definite matrix. It is known that every positive definite matrix $\mathbf{A}$ can be diagonalized as $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$, where
@@
is a $d$-dimensional diagonal matrix with $\lambda_{i}>0$ for i=1,2,,di=1,2,...,d, and $\mathbf{P}$ is an orthogonal matrix satisfying $\mathbf{P}^{-1}=\mathbf{P}^{\top}$. By expressing
@@
we can define the square root of $\mathbf{A}$ as $\mathbf{A}^{1/2}=\mathbf{P}\Lambda^{1/2}\mathbf{P}^{\top}$. It is then straightforward to verify that $(\mathbf{A}^{1/2})^{2}=\mathbf{A}$. $\mathbf{A}^{1/2}$ is also symmetric and positive definite.
Next, we will present some theorems related to the anisotropic Laplacian equation. While we believe these properties have been investigated in prior mathematical research, as they are closely related to the generalized Green’s function (Chen and Kuo, 2005), we are not aware of specific literature in the fields of cage or variational shape deformation that comprehensively addresses these results. Therefore, we provide a detailed discussion of them in the following.
@@
where $\delta(\mathbf{x})$ is the Dirac delta distribution satisfying $\delta(\mathbf{x})=0$ for $\mathbf{x}\neq\mathbf{0}$ and $\int_{\mathbb{R}^{d}}{\delta(\mathbf{x})=1}$, and $\omega_{d}={2\pi}^{d/2}/\Gamma(d/2)$ represents the surface area of the unit sphere in $\mathbb{R}^{d}$.
@@
We begin with the variable substitutions $\mathbf{x}=\mathbf{A}^{-1/2}\xi$ and $\mathbf{y}=\mathbf{A}^{-1/2}\eta$. According to Lemmas A.1 in the supplementary material, we have
@@
Since the Dirac delta distribution is a generalized function that violates the $\mathcal{C}^{2}$ condition of Lemma A.1, we introduce Lemma A.2 to establish the same variable substitution formula within the framework of distribution theory. Furthermore, within distribution theory, equality holds in the almost everywhere sense; thus, functions may differ on a set of measure zero. Nevertheless, for the fundamental solution, it is customary to seek a function that is smooth everywhere except at the singular point.
@@
The last equality follows from Lemma A.3 together with the identity $\det{\mathbf{A}^{1/2}}=|\det{\Lambda^{1/2}|}=\sqrt{\det{\mathbf{A}}}$. Therefore, we conclude that $\sqrt{\det{\mathbf{A}}}\ G_{\mathbf{A}}(\mathbf{x}-\mathbf{y})\$is the fundamental solution $G(\mathbf{u})$ of $d$-dimensional isotropic Laplacian equation, which is given by
@@
Moreover, the gradient $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)$ can be evaluated explicitly as:
(8) $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)=\begin{cases}\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-1}\mathbf{A}^{-1}(\xi-\eta)},&d=2,\\ \frac{1}{\omega_{d}\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-\frac{d}{2}}\mathbf{A}^{-1}(\xi-\eta)},&d\geq 3,\end{cases}$
will also be used frequently in the subsequent analysis.